OpenStudy (anonymous):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 32t + 384 shows the height f(t), in feet, of the sandbag after t seconds: Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. (4 points) Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? (4 points) Part C: Use your answer in part B to determine the axis of symmetry for f(x)? (2 points)

OpenStudy (anonymous):

@iambatman @mathmale @primeralph @Luigi0210

OpenStudy (primeralph):

Start.

OpenStudy (anonymous):

I have graphed it even though i need to explain why, so im just studying thegraph for help

OpenStudy (anonymous):

To begin you cvan remove a 16 from the entire equation

OpenStudy (anonymous):

Right?

OpenStudy (primeralph):

Just factor it.

OpenStudy (anonymous):

then you split the four into the two groups (-t^2-t)(-t+24)

OpenStudy (anonymous):

sorry 16(-t^2-t)(-t+24)

OpenStudy (anonymous):

set y to zero (-t^2-t)=0 and (-t+24)=0 then add t and to bothto get 24=t and -t^2=t

OpenStudy (anonymous):

wait what? @Kainui @iambatman

OpenStudy (anonymous):

@iambatman @mathslover @Compassionate

OpenStudy (primeralph):

Yeah, you have the procedure correct.